OpenStudy (anonymous):

how to find p if the distance between (p,3) and (-1,2p) is 5 units? ?

OpenStudy (amistre64):

could use vectors maybe?

OpenStudy (amistre64):

or just distance it lol

OpenStudy (anonymous):

but how?

OpenStudy (amistre64):

(p+1)^2 +(3-2p)^2 = 5^2

OpenStudy (amistre64):

p^2 +2p +1 +9 +4p^2 -12p = 25

OpenStudy (amistre64):

6p^2 -10p +10 - 25 = 0 6p^2 -10p -15 = 0 solve for p

OpenStudy (amistre64):

either complete the square method or quadratic formula it

OpenStudy (anonymous):

:) thx man... btw I feel awfully dumb after reading this... thx again

OpenStudy (amistre64):

:) anytime you wanna feel the fool; come on by :)